Problem 1: School A has a mean population IQ of 100 with a standard deviation of 15. A principal of the said school claims that the students in his school have above average intelligence. A random sample of 30 students IQ scores have a mean score of 112. Is there sufficient evidence to support the principal’s claim at 5% alpha? Problem: State the level of significance. *

1
Expert's answer
2022-02-17T04:01:18-0500

Solution:

We have that

"n=30"

"\bar x=112"

"\mu = 100"

"\sigma=15"

"H_0: \mu=100"

"H_a:\mu>100"

The hypothesis test is right-tailed.

The population standard deviation is known and the sample size is large (n≥30) so we use z-test.

Let the significance level be 5% in this test, therefore Z0.05 = 1.64

The critical region is Z > 1.64

Test statistic:



"Z_{test}=\frac{\bar x -\mu}{\frac{\sigma}{\sqrt n}}=\frac{112 -100}{\frac{15}{\sqrt {30}}}=4.38"

Since 4.38 > 1.64 thus the Ztest falls in the rejection region we reject the null hypothesis.

At the 5% significance level the data do provide sufficient evidence to support the claim. We are 95% confident to conclude that the students in the school are above average intelligence.